OpenStudy (anonymous):

Urgent! 10 minutes left to turn in! Identify the sequence graphed below and the average rate of change from n = 1 to n = 3. GRAPH IN NEXT POST

OpenStudy (anonymous):

an = 8(one half)n − 2; average rate of change is −6 an = 10(one half)n − 2; average rate of change is 6 an = 8(one half)n − 2; average rate of change is 6 an = 10(one half)n − 2; average rate of change is −6

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

@mathmale @mathmate

OpenStudy (anonymous):

@zepdrix, can you help? @mathmate, can you help?

zepdrix (zepdrix):

So the common ratio is 1/2. See how each term is getting cut in half each time? \[\Large\rm a_n=a_1\left(\frac{1}{2}\right)^{n-1}\]What about the first term?

zepdrix (zepdrix):

Oh they have a -2 in there >.< lemme fix that.

OpenStudy (anonymous):

Haha no worries!!

zepdrix (zepdrix):

I'll keep looking at it though, the -2 is throwing my brain off.

OpenStudy (anonymous):

That's okay, honestly, if you could just give me the equation I could probably figure it out.

zepdrix (zepdrix):

Ok ok ok I see what they did there. That was confusing. The -2 is part of the exponent.

zepdrix (zepdrix):

So we start with our general for the geometric sequence:\[\Large\rm a_n=a_1\left(\frac{1}{2}\right)^{n-1}\]

OpenStudy (anonymous):

You know what, you're right, I am so so sorry I wasn't more clear about that. I should have put ^n-2 or something!

zepdrix (zepdrix):

They're giving us coordinates in the form of: \(\Large\rm (n,~a_n)\) So we'll plug in the point \(\Large\rm (2,8)\) and solve for a_1 our first term.

zepdrix (zepdrix):

\[\Large\rm 8=a_1\left(\frac{1}{2}\right)^{2-1}\]

zepdrix (zepdrix):

So what do you get for your first term?

zepdrix (zepdrix):

Understand how to solve for \(\Large\rm a_1\) in that equation? :o

OpenStudy (anonymous):

Does a=16???

zepdrix (zepdrix):

Ok very good. Let's plug that into our function:\[\Large\rm a_n=16\left(\frac{1}{2}\right)^{n-1}\]So this is where things get weird. Maybe they wanted us to approach this differently.. I dunno. But anyway, see how the exponent is supposed to be n-2? So we need to take a 1/2 out of that whole thing.